OpenStudy (anonymous):

Describe an infinite geometric series with a beginning value of 2 that converges to 10.

OpenStudy (anonymous):

\[2+2r+2r^2+...=10\] \[\frac{2}{1-r}=10\] solve for \(r\)

OpenStudy (anonymous):

I got r = 4/5, I didn't know what to do after that though

OpenStudy (anonymous):

i think there is nothing to do after that

OpenStudy (anonymous):

\[\sum_{k=0}^{\infty}2\times \left(\frac{4}{5}\right)^k=10\]